# FREE-Merging: Fourier Transform for Model Merging with Lightweight Experts

### v1

@@ 4.3 Lightweight Expert Extraction
Since low-frequency signals are discarded in the backbone and information is lost during merging, this section constructs lightweight task experts to be added during inference, compensating for the task-specific capabilities. As analyzed in Sec  4.2, low-frequency signals contain task-specific information, but preserving them requires an inverse Fourier transform during each inference, which is inefficient. Thus, we propose a more efficient approach.
Since fine-tuning leads to task specialization, the most changed parameters during fine-tuning can be approximated as having the greatest impact on this process [27]. After reasonable rescaling, these selected parameters serve as task experts, minimizing storage (typically around 1% of parameters) and adding no extra computation during inference. The approximation’s validity is detailed in the Appendix 10.
Although previous work has proposed magnitude-based extraction [42], insufficient research on rescaling limits compression. Rescaling is crucial for expert extraction, so we propose a rescaling method. Let $M(v_{i},d)$ represent the top-$d$ parameters for task vector $v_{i}$, $\mu_{i}$ the rescale factor, and $e(v_{i})$ represents the expert for task $t_{i}$. Then we have:
@@
The goal is to maintain a consistent output after extraction while ensuring tasks with smaller backbone weights receive greater expert compensation to preserve performance.
In summary, we use an efficient method to extract task-specific experts, which can be integrated into the backbone to compensate for performance decline due to task conflicts without bring excessive storage and inference overhead.
4.4 FREE-Merging
Algorithm 1 Workflow of FREE-Merging
1:  FR-Merging: $\triangleright$ Only excute once.
3:  Expert Extraction: $\triangleright$ Only excute once.
4:  Get task experts $\{e_{i}\}_{i=1}^{K}$ as Eq. 4.
5:  Inference:
6:  for $x\in X$ do
7:     $[w_{1},...w_{K}]\leftarrow argmax(R(x))$
8:     $\theta_{*}=\theta_{m}+\sum_{i=1}^{K}w_{i}e_{i}$
9:     $Y\leftarrow Y\cup f(x;\theta_{*})$
10:  end for
10:  Output $Y$ for input $X$.
In this section, we describe how to combine the two parts into FREE-Merging, making it suitable for multiple tasks without affecting interference speed. Inspired by MoE, we use a router to dynamically assign tasks to different activated parameters. The router can take various forms [20], allowing for more flexibility in handling different tasks. For detailed algorithms, see Alg 1.
@@ 4.5 Time Complexity and Storage Space Analysis
In this section, we analyze the time complexity of merging and the storage space required after merging. We aim to keep both as small as possible to meet practical user needs. Assume that there are $n$ models to merge, each with $m$ parameters and requiring storage space $s$.
The time complexity analysis can be divided into two parts: merging and inference. In the merging process, the time complexity of FR-Merging is $O(nmlogm)$ [13] while that of extracting the expert is $O(nm)$. Thus, the overall time complexity of merging is $O(nmlogm)$, which is manageable. For example, merging three 10B models requires only $O(10^{12})$ addition operations, which is not a large load for current devices. Additionally, since only a lightweight router is added, the extra inference cost is negligible.
Next, we analyze storage space. We typically save only about 1% additional parameters per model. Therefore, our storage space requirement is $(1+0.01n)s$, which is a significant improvement compared to the requirement of $ns$ for the model ensemble. For large models, even a 1% reduction in data storage can greatly alleviate storage pressure. For detailed analysis, please refer to Sec 5.7.
@@ 5 Experiments
In this section, we evaluate the proposed method across various tasks. We compare FR-Merging with other cost-free methods (which do not require additional training or data), including Weight Averaging, Fisher Merging [28], Task Arithmetic [18], Ties-Merging [42], Breadcrumbs [6], and PCB-Merging [11]. We also compare FREE-Merging with other popular methods, including RegMean [19], AdaMerging [46], EMR-Merging [17], and Twin-Merging [24]. The original EMR-merging uses a perfect router; however, we use an imperfect router for realism. Some experiments include Traditional MTL with training data from all tasks as a baseline. The evaluation covers CV, NLP, and multi-modal tasks, followed by ablation studies and further analysis.
@@ 5.1 Merging Vision Models
Merging 8 ViTs. We select the two visual encoders of CLIP, ViT-B/32 and ViT-L/14 [32], as the pre-trained models and choose eight image classification tasks as target tasks: SUN397, Cars, RESISC45, EuroSAT, SVHN, GTSRB, MNIST, and DTD [46]. Additionally, an MLP is used as the lightweight router. Detailed information on the datasets and baselines can be found in the Appendix 9.1.
The comparison results of various methods can be found in Table 1. First, compared to existing methods that do not require additional training and data storage, our proposed FR-Merging shows significant improvements, achieving approximately 2.5% on ViT-B/32 and around 1.5% on ViT-L/14 over the current state-of-the-art, PCB-Merging [11]. This indicates that the Fourier filtering effectively removes task-specific parameters from the task vector, greatly enhancing the generalization ability of the model and reducing performance loss due to task conflicts. Through FR-Merging, we can effectively construct a merged backbone that can be used independently at a low cost. Second, compared to methods requiring extra information, FREE-Merging also shows a considerable improvement about 2%. This advantage is based on the lightweight computation and efficient expert extraction of FREE-Merging, as we achieved this with 1% additional storage for each model, outperforming 3% of EMR-Merging [17] and 2% of Twin-Merging [24]. Thus, FREE-Merging provides an excellent balance between storage, computation, and performance.
@@
Merging 30 ViTs. We select ViT-B/16 as the pre-trained model and evaluate the model merging performance across 30 image classification tasks. The dataset selection and configuration of fine-tuning tasks follow [17].
The results are shown in Table 2. Details can be found in Appendix 11.1. It can be found that, compared to current methods, our proposed FR-Merging and FREE-Merging show remarkable performance. Although real applications rarely involve merging such a large number of models, we introduce this experiment mainly to demonstrate that the proposed Fourier filtering and lightweight task experts can mitigate performance loss caused by task conflicts even in extreme scenarios. This further illustrates the effectiveness and broad applicability of our methods.
5.2 Merging Language Models
Merging Medium-sized Language Models. We investigate the generalization ability of our method on medium-scale language models, using RoBERTa [50] as the pre-trained model with full fine-tuning on eight discriminative tasks [42]. Details can be found in Appendix 9.1.
The average results of the eight tasks after merging are shown in Table 3, and the complete results are available in Appendix 11. It can be observed that among our proposed methods, FR-Merging shows a significant advantage of nearly 2% over other cost-free methods. Additionally, FREE-Merging also exhibits remarkable effectiveness, achieving optimal results on almost all datasets and outperforming current methods in average performance. This demonstrates the applicability and great potential of our proposed methods for language models.
Firstly, we use T0-3B [33] as the pre-trained model and $(IA)^{3}$ [22] as the PEFT method, fine-tuning on 11 discriminative tasks [17]. Our proposed FR-Merging and FREE-Merging show great advantages with an improvement of nearly 1% in Table 3. Then, we select Qwen-14B [2] as the pre-trained model and LoRA [16] as the PEFT way, fine-tuning on three generative tasks before merging [24]. The results show that our method exhibits nearly a 2% advantage, whether in cost-free conditions or the complete framework. This demonstrates that our methods remain highly feasible even with PEFT, confirming their generalization ability. Full results are available in the Appendix 11.
@@ 5.2 Merging Language Models
Merging Large Language models. We investigate the merging of large models. We use LLaMa2-13B [35] as the pre-trained model, utilizing WizardLM [41] for instruction following, WizardMath [25] for math, and WizardCoder-Python [26] for coding. Experiments are conducted on the AlpacaEval (instruction following task), GSM8K (math task), and MBPP (code generation task) [47].
The results are shown in Table 4. It can be observed that our proposed FR-Merging even surpasses existing methods that rely on additional information, showing great advantages. This indicates that our findings in FR-Merging regarding low-frequency signals and task generalization also hold when extended to large models, which is highly significant for the practical deployment of large models. Additionally, our proposed FREE-Merging also shows considerable advantages. This suggests that FREE-Merging can also be effectively applied to large models, promoting rapid deployment of large models across cross-domain tasks.
@@ 5.3 Merging Multi-Modal Models
We merge various fine-tuned models of the BEiT3-base [37] across five different tasks: ImageNet-1k (Image Classification), VQAv2 (Visual Question Answering), NLVR2 (Visual Reasoning), COCO Captioning (Image Captioning), and COCO Retrieval (Image-Text Retrieval). COCO Captioning is evaluated using BLEU4, CIDEr, METEOR, and ROUGE-L, while the other tasks are measured by accuracy [17]. Details are available in Appendix 9.1.
The results in Table 5 demonstrate that our method also shows significant improvements in multi-modal models. Unlike the merging of vision or language models, which often involve similar tasks, such as classification or generation, the task similarity here is relatively low. Therefore, the superiority in this configuration fully indicates the generalization ability of our FR-Merging and FREE-Merging, making them suitable for a wide range of applications.
Figure 3: Visualization of the loss landscape on EuroSAT before and after high-pass filtering (top vs bottom) in the model merging of the fine-tuned ViT-B/32 on EuroSAT and DTD tasks [18].
@@ 5.4 Fourier Transform Analysis
In this section, we explain why Fourier high-pass filtering improves merging performance from an optimization perspective. Previous work [21] shows that landscape geometry impacts model generalization. Fig 3 demonstrates that high-pass filtering widens the valley in the loss landscape (bottom plot). Since fine-tuned models from the same pre-trained model are closer in the landscape [30], a wider valley increases the likelihood of merged results falling within the loss basin as shown in the bottom plot of Fig 3, leading to better performance. Without filtering, the landscape is flat at a high level, which has little effect on model merging and only highlights the advantage of pre-training [15].
Ablation Part Avg. Acc(%)
Backbone Weight Averaging 65.80
+ FR-Merging 76.96 ($\uparrow$11.16)
+ $\lambda$ as Equa. 3 77.64 ($\uparrow$11.84)
Expert + Top-K Expert 79.38 ($\uparrow$13.58)
+ Our Rescaling 89.68 ($\uparrow$23.88)
Table 6: Ablation study of FREE-Merging. Gradually add each module proposed in this paper from top to bottom.
@@ 5.5 Ablation Study
Ablation on FR-Merging. In this part, we analyze the impact of replacing FR-Merging on the backbone performance and overall performance. Table 6 shows that adding FR-Merging leads to an 11.16% performance increase. Next, we replace high-pass filtering with other cost-free methods such as Ties [42], Top-K, Task Arithmetic [18], outlier dropout [6], EMR-Merging [17], and PCB-Merging [11], while also comparing low-pass filtering, and median filtering, with detailed results in Fig 4(a). To ensure a fair comparison, all experts use an additional 1% of parameters.
@@
Figure 4: a) The ablation experiments on FR-Merging show that our proposed method achieves optimal performance. b) Performance of experts at different sparsity rates. c) In experiments with random routers, our approach demonstrates optimal performance.
Our proposed FR-Merging outperforms other cost-free methods on both the backbone and in combination with task experts, demonstrating its effectiveness in minimizing task conflicts in the merged network. Examining the different filtering methods, we find that low-pass filtering leads to great performance degradation, confirming that the low-frequency information contains task-specific signals and greatly affects generalization. Median-pass filtering also performs worse than FR-Merging, and this is because some high-frequency signals preserve abrupt changes in the signal and their removal harms the performance of the model.
Ablation on Merging Coefficients. This part presents an ablation study on our merging coefficients as in Eq. 3. Table 6 shows the results. It can be found that incorporating merging coefficients instead of simply averaging leads to a slight improvement. This is because the merging coefficients fully consider the variations of the outputs for each task before and after merging, ensuring that the output after merging is similar to the one before merging. This is especially suitable for cases where the magnitudes of the fine-tuning parameters of different tasks are inconsistent.
Ablation on Task Expert. We conduct an ablation study on the expert extraction method, First, Table 6 shows that incorporating our proposed experts and rescaling method leads to great improvements, confirming the effectiveness.
To validate the superiority of our method, we evaluate the performance of only utilizing experts while retaining different proportions of parameters. We compare the original Top-K selection, Bernoulli selection [47], SVD extraction [24], and our proposed Top-K extraction with rescaling. As shown in Fig 4(b), our method shows substantial advantages, particularly under high sparsity conditions, supporting the effectiveness of our lightweight expert extraction.
5.6 Router Analysis
In this part, we analyze the proposed router method. We aim to achieve strong performance even if the router performs sub-optimally. Router inaccuracies are likely to occur due to insufficient training resources or issues during the training on edge deployment. Therefore, the advantages of our methods in this scenario are of great practical significance.
We use a random router to simulate an inaccurate router and compare our method with several existing router-based methods. Our proposed method outperforms Twin-Merging by 13.5% and outperforms EMR-Merging by 8.5% on average, as shown in Fig 4(c), demonstrating that our method constructs a reasonable backbone and generalizable experts.
@@ 5.7 Speed and Storage Space Analysis
This part analyzes the training costs, inference speed, and storage requirements; details are shown in Table 7. Our FR-Merging achieves strong performance with a single model for storage, significantly reducing storage demands without extra training or inference costs. FREE-Merging, with a small increase in training and storage overhead, further enhances performance. These two methods greatly facilitate the practical deployment of deep learning at minimal cost.

### v2

@@ 5.3 Lightweight Expert Extraction
To compensate for information lost due to discarding low-frequency signals, we utilize lightweight task experts added during inference. As analyzed in Sec. 5.2, low-frequency signals contain task-specific information, but preserving them requires an inefficient inverse Fourier transform during each inference. Thus, we propose a more efficient way.
Since fine-tuning leads to task specialization, the most changed parameters can be approximated as having the greatest impact [32]. After reasonable rescaling, these selected parameters serve as task experts, minimizing storage (typically around 1% of parameters) without extra inference computation. The validity is detailed in Appendix D.4.
Although magnitude-based extraction [48] exists, limited research on rescaling hinders compression. As rescaling is crucial to expert extraction, we propose a rescaling method. Let $M(v_{i},d)$ represent the top-$d$ ($d$ is a percentage) parameters for task vector $v_{i}$, $\mu_{i}$ be the rescale factor, and $e(v_{i})$ be the expert for task $t_{i}$. Then we have:
@@
The goal is to maintain a consistent output after extraction to preserve performance. Details are in Appendix D.5.
In summary, we efficiently extract task-specific experts to integrate into the backbone, mitigating task conflicts without excessive storage or inference overhead.
@@ B L B L B L B L B L B L B L B L B L
5.4 FREE-Merging
Algorithm 1 Workflow of FREE-Merging
1:  FR-Merging: $\triangleright$ Only excute once.
3:  Expert Extraction: $\triangleright$ Only excute once.
4:  Get task experts $\{e_{i}\}_{i=1}^{K}$ as Eq. 4.
5:  Inference:
6:  for $x\in X$ do
7:     $[w_{1},...w_{K}]\leftarrow argmax(R(x))$
8:     $\theta_{*}=\theta_{m}+\sum_{i=1}^{K}w_{i}e_{i}$
9:     $Y\leftarrow Y\cup f(x;\theta_{*})$
10:  end for
10:  Output $Y$ for input $X$.
This section presents FREE-Merging, combining both parts to make it suitable for multiple tasks without affecting inference speed. Inspired by MoE [39], a router dynamically assigns tasks to activated experts. The router can take various forms [24], allowing for more flexibility in handling different tasks. Detailed discussions about router can be found in Appendix D.6. For detailed algorithms, see Alg 1.
@@ 5.5 Time Complexity and Storage Space Analysis
This section analyzes the merging time complexity and storage requirements. We aim to minimize both for practical use. Assume there are $n$ models to merge, each with $m$ parameters and requiring storage space $s$.
The time complexity includes the merging and inference part. In the merging part, FR-Merging has a complexity of $O(nmlogm)$ [14] while expert extraction is $O(nm)$, making the overall merging complexity $O(nmlogm)$, which is manageable. For instance, merging three 10B models requires only $O(10^{12})$ addition operations, a minor load for modern devices. Moreover, with only a lightweight router added, the extra inference cost is negligible.
Next, we analyze storage cost. We typically save only about 1% additional parameters per model, requiring $(1+0.01n)s$ storage, which is far less than the requirement of $ns$ for the model ensemble. For large models, reducing storage requirements is crucial for alleviating storage pressure on edge users. For detailed analysis, please refer to Sec 6.7.
@@ 6 Experiments
In this section, we evaluate the proposed methods across various tasks. We compare FR-Merging with other cost-free methods (which do not require additional training or data), including Weight Averaging, Fisher Merging [33], Task Arithmetic [21], Ties-Merging [48], Breadcrumbs [7], and PCB-Merging [12]. We also compare our FREE-Merging with popular methods, including RegMean [23], AdaMerging [52], EMR-Merging [20], and Twin-Merging [29]. The original EMR-Merging uses a perfect router; however, we use an imperfect router for realism. Some experiments include traditional MTL with data from all tasks as a baseline. The evaluation covers CV, NLP, and multi-modal tasks.
@@ 6.1 Merging Vision Models
Merging 8 ViTs. We select the two visual encoders of CLIP, ViT-B/32 and ViT-L/14 [37], as the pre-trained models and choose eight image classification tasks as target tasks: SUN397, Cars, RESISC45, EuroSAT, SVHN, GTSRB, MNIST, and DTD [52]. Additionally, an MLP is used as the lightweight router. Detailed information on the datasets and baselines can be found in the Appendix C.1.
@@
Table 2 presents the comparative results. First, FR-Merging outperforms other cost-free methods, improving ViT-B/32 by 2.5% and ViT-L/14 by 1.5% over PCB-Merging [12], demonstrating its effectiveness in reducing performance loss caused by task interference. Through FR-Merging, we can construct a merged backbone that can be used independently at a low cost. Here, we focus on the basic FR-Merging. In Sec. 6.6, we discuss further improvements achieved by combining it with existing methods. Then, compared to methods requiring extra information, FREE-Merging also improves performance by 2%, leveraging lightweight computation and efficient expert extraction, requiring only 1% additional storage per model, outperforming 3% of EMR-Merging [20] and 2% of Twin-Merging [29]. Thus, FREE-Merging provides an excellent balance between storage, computation, and performance.
Merging 30 ViTs. We select ViT-B/16 as the pre-trained model and evaluate the model merging performance across 30 image classification tasks following [20].
Table 3 presents the results, with detailed analysis provided in Appendix E.1. It is evident that our proposed FR-Merging and FREE-Merging exhibit remarkable performance. Although merging a large number of models is uncommon in practice, this experiment underscores the effectiveness of Fourier filtering and lightweight task experts in mitigating performance loss caused by task interference. These results highlight the robustness and broad applicability of our methods across diverse scenarios.
6.2 Merging Language Models
Merging Medium-sized Language Models. We investigate the generalization ability of our method on medium-scale language models, using RoBERTa [58] as the pre-trained model with full fine-tuning on eight discriminative tasks [48]. Details can be found in Appendix C.1.
The average merging results of the eight tasks are presented in Table 4, with complete results in Appendix E.2. It can be observed that, FR-Merging outperforms other cost-free methods by nearly 2%. Additionally, FREE-Merging also exhibits remarkable effectiveness, achieving optimal results on almost all datasets and surpassing current methods in average performance. This highlights the great potential of our proposed methods for language models.
Firstly, we use T0-3B [38] as the pre-trained model and $(IA)^{3}$ [26] as the PEFT method, fine-tuning on 11 discriminative tasks [20]. Our proposed FR-Merging and FREE-Merging improve performance by nearly 1%, as shown in Table 4. Next, we employ Qwen-14B [2] with LoRA [19] as the PEFT way, fine-tuning on three generative tasks [29]. The results show that our methods lead to a 2% advantage in both cost-free and full frameworks. This confirms the feasibility with PEFT and strong generalization of our methods. Full results are available in the Appendix E.2.
@@ 6.2 Merging Language Models
Merging Large Language models. We explore large model merging using LLaMa2-13B [40] as the pre-trained model, with WizardLM [46] for instruction following, WizardMath [30] for math, and WizardCoder-Python [31] for coding. Experiments are conducted on AlpacaEval (instruction following), GSM8K (math), and MBPP (code) [53].
Table 5 shows that our proposed FR-Merging even surpasses existing methods that rely on additional information. This indicates that our findings in FR-Merging regarding low-frequency signals and task interference also hold when extended to large models. This is crucial for practical usage of large models. Additionally, our proposed FREE-Merging also demonstrates strong advantages, enabling efficient cross-domain deployment of large models.
@@ 6.3 Merging Multi-Modal Models
We merge various fine-tuned BEiT3-base [42] models across five tasks: ImageNet-1k (Classification), VQAv2 (Visual Question Answering), NLVR2 (Reasoning), COCO Captioning (Image Captioning), and COCO Retrieval (Image-Text Retrieval). COCO Captioning is evaluated with BLEU4, CIDEr, METEOR, and ROUGE-L, while the others use accuracy [20]. Details are in Appendix C.1.
Table 6 shows that our methods gain great improvements in multi-modal models. Unlike vision or language model merging, which often involve similar tasks, such as classification, the task similarity here is relatively low. The strong performance highlights the generalization ability of our methods, making them widely applicable.
Figure 4: Visualization of the loss landscape on EuroSAT before and after high-pass filtering (top vs. bottom) in the model merging of the fine-tuned ViT-B/32 on EuroSAT and DTD tasks [21].
@@ 6.4 Fourier Transform Analysis
In this section, we explain why Fourier high-pass filtering improves merging performance from an optimization perspective. Previous work [25] shows that landscape geometry impacts model generalization. Fig. 4 demonstrates that high-pass filtering widens the valley in the loss landscape (bottom plot). Since fine-tuned models from the same pre-trained model are closer in the landscape [34], a wider valley increases the probability of merged results falling within the loss basin as shown in the bottom plot of Fig. 4, leading to better performance. Without filtering, the landscape is flat at a high level, which has little effect on model merging and only highlights the advantage of pre-training [17].
FR-Merging Equa.3 Top-K Expert Scaling ViT-B/32 ViT-B/16 ViT-L/14
70.10 75.24 84.50
76.96 81.34 88.21
71.43 77.41 84.20
78.10 82.54 88.30
65.24 68.98 70.13
79.38 84.31 89.53
82.32 86.87 90.47
89.68 91.09 93.70
Table 7: Ablation of FREE-Merging components on 8 visual tasks.
@@ 6.5 Ablation Study
Ablation on FR-Merging. We analyze the impact of FR-Merging on the backbone and overall performance. Table 7 shows that the performance of various models greatly improves after employing FR-Merging. We then compare high-pass filtering with other cost-free methods, including Ties [48], Top-K, Task Arithmetic [21], outlier dropout [7], EMR-Merging [20], and PCB-Merging [12], as well as low-pass and band-pass filtering, with results in Fig. 5(a). For fairness, all experts use 1% additional parameters.
@@
Figure 5: a) The ablation experiments on FR-Merging show that our proposed methods achieve optimal performance. b) Performance of experts at different sparsity rates. c) In experiments with random routers, our approach demonstrates optimal performance.
Method FR-Merging ViT-B/32 ViT-L/14
Ties-Merging [48] 73.6 86.0
Ties-Merging [48] 78.3 88.4
Breadcrumbs [7] 75.0 86.4
Breadcrumbs [7] 78.8 88.9
TSVM [16] 83.9 91.9
TSVM [16] 84.8 92.5
Table 8: Transferability analysis of FR-Merging on 8 visual tasks.
Our proposed FR-Merging outperforms other cost-free methods on both the backbone and with task experts, minimizing task interference. Examining the different filtering methods, we find that low-pass filtering leads to great performance degradation, confirming that task interference in low-frequency regions is severe. Band-pass filtering also performs worse than FR-Merging, and this is because some high-frequency signals preserve abrupt changes in the signal and their removal harms the performance of the model.
Ablation on Merging Coefficients. We conduct an ablation study on our merging coefficients as in Eq. 3, with results in Table 7. Using them instead of simple averaging yields slight improvements by considering output variations, benefiting tasks with varying fine-tuning magnitudes.
Ablation on Task Expert. We conduct an ablation study on the expert extraction way. Table 7 shows that incorporating our experts and rescaling method results in significant improvements, confirming their effectiveness. To validate the superiority of our method, we evaluate expert-only performance with varying parameter proportions, comparing Top-K selection, Bernoulli selection [53], SVD extraction [29], and our Top-K selection with rescaling. Fig 5(b) shows that our method excels, particularly under high sparsity.
Router Analysis. We analyze the proposed router method, aiming for strong merging performance despite sub-optimal routing. Since router inaccuracies may arise from limited training resources or edge deployment issues, our methods maintain performance in real-world scenarios. Using a random router to simulate inaccuracy, we compare our method with existing router-based approaches. Our method outperforms Twin-Merging by 13.5% and EMR-Merging by 8.5% on average, as shown in Fig 5(c), indicating that our methods construct a reasonable backbone and generalizable experts.
6.6 Transferability Analysis
Our proposed FR-Merging can be seen as a method that can integrate with others, as it addresses frequency-domain task interference that current methods overlook. As shown in Table 8, combining various methods with FR-Merging leads to performance improvements. Therefore, FR-Merging can be integrated with more approaches to achieve greater benefits.
@@ 6.7 Speed and Storage Space Analysis
This part analyzes training costs, inference speed, and storage as shown in Table 9. FR-Merging achieves strong performance with significantly reducing storage demands without extra training or inference costs, while FREE-Merging slightly increases overhead for better performance. Both of these two methods enhance deep learning deployment.